OpenStudy (anonymous):

Using the function find f(g(x)) f(x)= 7x-3 g(x)=x^2+5

OpenStudy (jdoe0001):

\(\bf f(x)= 7x-3\qquad \qquad g(x)=\color{red}{x^2+5}\\ \quad \\ f(\quad g(x)\quad ) = 7(\color{red}{x^2+5})-3\)

OpenStudy (anonymous):

How you solve for it is you take the expression for g(x) and plug that it as the value for x in the f(x) function.

OpenStudy (shamil98):

ooh , my bad , i f(g(x)) = (f o g)(x)

OpenStudy (anonymous):

So it would be 7(X^2+5-3) and i solve that?

OpenStudy (shamil98):

F(g(x)) = 7(x^2 +5) -3 = 7x^2 + 35 -3 = 7x^2 + 32

OpenStudy (shamil98):

No, you don't solve for that, that is the answer. ^

OpenStudy (anonymous):

Alright thank you!